OpenStudy (anonymous):

How many grams of a stock solution that is 87.5 percent H2SO4 by mass would be needed to make 275 grams of a 55.0 percent by mass solution? Show all of the work needed to solve this problem.

OpenStudy (anonymous):

Your final solution contains 55% by mass of H2SO4. You have 275 g of that solution, so it contains 0.55 x 275 g of H2SO4 i.e. 151.25 g H2SO4. The stock solution contains 87.5% H2SO4. So 100 g of that solution contains 87.5 g of H2SO4. You need 151.25 g of H2SO4, so now it’s just a simple proportion question: if you call the mass of stock solution m, then … m/100 = 151.25 / 87.5 so m = 100 (151.25 / 87.5) = 173.8 g